The drill pipe was raised 4 m and Hole 1010B was spudded at 1945 hr on 25 April. The water depth was established at 3475.3 meters below rig floor (mbrf) based on recovery of the mudline core. APC Cores 167-1010B-1H through 3H were taken from 0 to 23.2 mbsf with 102.2% recovery. A sudden power failure in the auxiliary 480-V transformer temporarily caused a generator to drop offline. The drill pipe was pulled above the seafloor, ending Hole 1010B while repairs were made to the generator.
The vessel was offset 10 m to the north and Hole 1010C was spudded at 0200 hr on 26 April. The water depth was established at 3476.5 mbrf based on recovery of the mudline core. APC Cores 167-1010C-1H through 17H were taken from 0 to 157.5 mbsf with 103.9% recovery. Oriented cores were obtained starting with Core 3H. XCB Cores 18X through 23X were taken down to 209.4 mbsf with 55% recovery. Excellent recovery was obtained on XCB Cores 18X through 20X, but a chert layer found on Core 21X broke off several of the tungsten carbide inserts on the XCB cutting shoe, resulting in a jammed core liner. The inserts were recovered in the core catcher. XCB Cores 22 through 23X were taken with a carbonado diamond shoe. Core 22X suffered from core jamming, and Core 23X penetrated basalt, and was terminated after 5.6 m advancement in 45 min. One MDCB core was taken down to 213.9 mbsf, and a solid section of basalt was recovered (cored 4.5 m, recovered 0.91 m).
Hole 1010D was spudded at 0600 hr on 27 April. The water depth was established at 3477.5 mbrf. APC Cores 167-1010D-1H through 6H were taken from 0 to 51.5 mbsf with 106% recovery. Adara heat-flow measurements were performed on Cores 3H through 6H. Following Core 6H, the coring wireline became wrapped around the sinker bar assembly, trapping the core barrel in the BHA. Attempts to free the core barrel were unsuccessful. The Kinley cutter was dropped, and the coring line was sheared immediately above the core barrel allowing the coring line to be retrieved. The drill string was pulled clear of the seafloor at 1645 hr on 27 April.
After pulling the drill pipe and bottom-hole assembly (BHA) to the surface, the core barrel was dislodged from the BHA, and the pipe was tripped back to bottom. Hole 1010E was spudded at 0430 on 28 April. The water depth was established at 3476.5 mbrf. APC Cores 167-1010E-1H through 16H were taken from 0 from 151.5 mbsf with 103.1% recovery). XCB Cores 17X to 19X were taken to 180.2 mbsf with 100.1% recovery. The pipe was pulled clear of the seafloor at 2200 hr 28 April, ending Hole 1010E.
The vessel was offset 10 m to the south and Hole 1010F was spudded at 2300 hr 28 April. The water depth was established at 3476.3 mbrf. APC Cores 1H through 6H were taken down to 55.7 mbsf with 102.3% recovery. The drill string was tripped back to the surface and secured for the 8-hr transit to Site 1011 by 0945 29 April.
Hole 1011A was spudded, and a single APC core (Core 167-1011A-1H) was taken from 0 to 9.5 mbsf. A full barrel prevented the establishment of an accurate mudline, and the hole was abandoned.
The drill pipe was raised 3 m, and Hole 1011B was spudded at 2245 hr on 29 April. The water depth was established at 2032.5 mbrf based on recovery of the mudline core. APC Cores 1671011B-1H through 15H were taken from 0 to 137.9 mbsf with 103.8% recovery. XCB Cores 16X through 31X were taken down to 281.5 mbsf with 89.1% recovery. Hole 1011B was terminated after recovery of several sections of vesicular basalt. A wiper trip was made from 281.5 mbsf up to 78 mbsf and back to the bottom of the hole where 3 m of fill was found. The pipe was raised to 78.5 mbsf and preparations for logging were made. Hole 1011B was logged with the Triple Combination tool string (density, neutron porosity, resistivity, and natural gamma ray) from 2310 to 2113 mbrf (277.5 to 79.5 mbsf). The FMS/Sonic/Gamma Ray (GR) tool string was run from 2310 to 2128 mbrf (277 to 95 mbsf). Both logging runs obtained excellent results. The pipe was pulled clear of the seafloor at 1830 hr 1 May, ending Hole 1011B.
Hole 1011C was spudded at 1900 hr on 1 May. APC Cores 167-1011C-1H through 15H were taken from 0 to 136.6 mbsf with 103.8% recovery. XCB Cores 1011C-16X through 20X were taken down to 184.2 mbsf with 101.2% recovery. The pipe was pulled clear of the seafloor at 1215 hr 2 May, ending Hole 1011C.
Hole 1011D was spudded at 1245 hr on 2 May. APC Cores 167-1011D-1H and 2H were taken from 0 to 16.9 mbsf with 102% recovery. The pipe was pulled clear of the seafloor at 1330 hr 2 May, ending Hole 1011D.
Hole 1011E was spudded at 1345 on 2 May. APC Cores 1011E-1H through 16H were taken down to 142.3 mbsf with 104.4% recovery. The drill string was tripped back to the surface and cleared the rotary table at 0130 hr 3 May, ending Hole 1011E. The JOIDES Resolution was underway for the 7.5-hr transit to Site 1012 by 0145 hr 3 May.
SITE 1012 (PROPOSED SITE BA-1)
Transit from Site 1011 to Site 1012
The 69.0-nmi transit from Site 1011 to Site 1012 (proposed Site BA-1) was accomplished in 5.75 hr at an average speed of 11.5 kt. A 3.5-kHz PDR survey was performed while approaching Site 1012. A Datasonics 354M beacon was dropped on GPS coordinates at 0800 hr on 3 May.
Hole 1012A was spudded at 1145 hr on 3 May. APC Cores 167-1012A-1H through 13H were taken from 0 to 118.7 mbsf with 105.7% recovery. Oriented cores were obtained starting with Core 3H. XCB Cores 14X through 30X were taken to 273.5 mbsf with 93.2% recovery. Headspace methane concentration increased with increasing depth, but no significant higher weight molecular hydrocarbons other than C1 and C2 were observed, indicating that the methane is of biogenic origin and not significant for safety and pollution investigations. The hole was displaced with heavy mud at the completion of coring operations.
Hole1012B was spudded at 1530 hr on 4 May. APC Cores 167-1012B-1H through 14H were taken from 0 to 132.3 mbsf with 99.2% recovery. Adara temperature measurements were taken on Cores 4H, 6H, 8H, and 10H.
Hole 1012C was spudded at 0115 hr on 5 May. APC Cores 167-1012C-1H through 9H were taken down to 82.6 mbsf with 105.2% recovery. The drill string was tripped back to the surface and secured for the 3.25-hr transit to Site 1013 by 0945 hr on 5 May.
Hole 1013A was spudded at 1730 hr on 5 May. APC Cores 167-1013A-1H through 10H were taken from 0 to 90.6 mbsf with 100.8% recovery. Oriented cores were obtained starting with Core 3H. XCB Cores 11X through 16X were taken to 146.1 mbsf with 82.2% recovery.
Hole 1013B was spudded at 0730 hr on 6 May. APC Cores 167-1013B-1H through 10H were taken from 0 to 94.1 mbsf with 104.8% recovery. XCB Core 167-1013B-11X was taken down to 103.8 mbsf with 96.8% recovery. Hole 1013B was originally scheduled to penetrate to 150 mbsf, but was cut short because of the constraint of having to arrive in San Diego at 0500 hr on 7 May.
Hole 1013C was spudded at 1430 hr 6 May. APC Cores 167-1013C-1H through 8H were taken down to 72.2 mbsf with 102.4% recovery. Adara temperature measurements were taken on Cores 4H, 6H, and 8H. The time for Site 1013 expired and the drill string was tripped back to the surface and secured for the 8-hr transit to San Diego by 2115 hr 6 May.
SITE 1014 (PROPOSED SITE CA-15A)
Transit from San Diego to Site 1014
The 151.0-nmi transit from San Diego to Site 1014 (proposed Site CA-15A) was accomplished in 15.5 hr at an average speed of 10.4 kt. A 3.5-kHz PDR survey was performed while approaching the site. The JOIDES Resolution arrived at Site 1014 at 0530 hr on 8 May.
Hole 1014A was spudded at 0900 hr on 8 May. APC Cores 167-1014A-1H through 6H were taken from 0 to 50.6 mbsf with 100.8% recovery. Oriented cores were obtained starting with Core 3H. APC refusal was reached at 50.6 mbsf when 70,000 lb of overpull was required to free the barrel. XCB Cores 7X through 50X were taken down to 449.0 mbsf with 88.7% recovery. Hole 1014A was logged with the Triple Combination, the FMS/Sonic/GR, and the GHMT tool strings from 443 to 58 mbsf with good results. The vessel was offset 100 m to the east in an attempt to find an area more conducive to APC coring.
Hole 1014B was spudded at 1915 hr on 10 May. APC Cores 167-1014B-1H through 13H were taken from 0 to 114.7 mbsf with 100.8% recovery. Apparently, the 100-m offset was enough to allow the piston coring to continue deeper than at Hole1014A before APC refusal. Adara temperature measurements were taken on Cores 3H through 6H and on Core 8H, and three good data points were acquired. While pulling the core liner from the core barrel following Core 13H, the first meter of the liner exploded. The remaining liner was still inside the core barrel and shattered, but remained intact enough to be pulled out of the core barrel. The portion of the liner that exploded caused part of the core to fall on the rig floor. XCB coring continued with Cores 167-1014B-14X through 27X, and was taken down to 245.0 mbsf with 83.6% recovery.
Hole 1014C was spudded at 1315 hr on 11 May. APC Cores 167-1014C-1H and 2H were taken down to 19.5 mbsf with 102.8% recovery.
Hole 1014D was spudded at 1400 hr on 11 May. APC Cores 167-1014D-1H through 10H were taken down to 92.0 mbsf with 104.5% recovery. XCB Cores 167-1014D-11X through 13X were taken down to 120.9 mbsf with 83.8% recovery. The drill string was tripped back to the surface and secured for the 8-hr transit to Site 1015 by 2145 hr on 11 May.
Hole 1015A was spudded at 0815 hr on 12 May. APC Cores 167-1015A-1H through 16H were taken from 0 to 149.5 mbsf with 93.1% recovery. Oriented cores were obtained on Cores 3H through 9H. An Adara temperature measurement was run on Core 4H. Further Adara measurements were canceled because of overpull and the slow drilling advancement. Advancement of the APC bit was very slow through the entire APC sequence. The recovered core exhibited multiple turbidites, which contributed to the slow drilling conditions. The vessel was offset 100 m to the east in an attempt to find an area more conducive to APC coring.
Hole 1015B was spudded at 2000 hr on 12 May. APC Cores 167-1015B-1H through 12H were taken from 0 to 97.8 mbsf with 83.2% recovery. The APC coring system encountered refusal in Core 12H, and the hole was terminated. The drill string was tripped back to the surface and secured for the 16-hr transit to Site 1016 by 0545 hr on 13 May.
Site 1016-1022 Operations
To 167 Table of Contents